Using substitution to eliminate one of the variables. Show your work
x+y=7
2x+3y=21

Answer by TutorBecky(6)

x+y=7
2x+3y=21
Work::
x + y = 7
-x -x
----------
y = -1x +7 :::: Just so you don't get confused x is the same as 1x
Plug the equation into the other for y
2x + 3 (1x +7 ) = 21 Distribute 3*1x+3*7
2x + 3x +21 = 21 Combine the same variables : 2x+3x
5x + 21 = 21
- 21 -21 Subtract 21 on both sides
------------
5x =0 Divide by 5 on both sides
-----
5 5
x = 0
Now to get y
x + y = 7
0 + y = 7
y =7

Now to check if your right;; Plug in your varibles
x+y=7
0+7= 7
7 = 7 Check!
------------
2(0) + 3 (7) = 21
0 +21 = 21
21=21 Check!
Hope i helped